OpenStudy (supercait10):

What type of number is shown below? 0.313311333111... A. Repeating decimal B. rational C. irrational D. terminating decimal

ganeshie8 (ganeshie8):

what are your first thoughts ? do you see any obvious options that can be eliminated ?

OpenStudy (supercait10):

Select all the rational numbers from the list. a \[\sqrt{2}\] b. 8.212211222111... c. 8.333333333 d \[\sqrt{80}\] e. \[\sqrt{}\] Select all the irrational numbers from the list. a. \[\sqrt{}\] b. \[\sqrt{}\] c. 813.1 d. -9 e. 8.070070007... Which of the following numbers is irrational? a. \[\sqrt{36}\] b. \[\frac{ 1 }{ 5 }\] c. \[\sqrt{60}\] d. 6.3

OpenStudy (supercait10):

rational

OpenStudy (mathmath333):

this is an irrational number i think